Find all real solutions of the equation.
x^2 + 5x - 6 = 0

Answer by ewatrrr(24785)

 
Hi,
x^2 + 5x - 6 = 0
(x + 6)(x -1) = 0 x = -6 or x = 1
Use FOIL to check the factoring
F First terms x^2
O Outside terms -x
I Inside terms +6x
L Last terms -6